OpenStudy (anonymous):

278^4x=37,800 find x

OpenStudy (lgbasallote):

\[\huge 278^{4x} = 37,800\] is that the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

anyone know?

OpenStudy (anonymous):

Take Log base 278 on both sides so you get 4x = Log base 278 of 37800 Then on right hand side use change of base formula to change it to Ln 37800/Ln 278 so u get 4x = Ln 37800/Ln 278 solve right side on calculator and divide that answer by 4 to get x

OpenStudy (anonymous):

i tried and i dont get the correct answer!

OpenStudy (anonymous):

i need the answer please

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

I am getting x = 0.47

OpenStudy (anonymous):

i got that too and its not correct thanks anyways though

OpenStudy (anonymous):

whats the correct answer ?

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

\[4x=\frac{\ln(37,800)}{\ln(278)}\] \[x=\frac{4\ln(37,800)}{\ln(278)}\] i get 7.49 http://www.wolframalpha.com/input/?i=4ln%2837800%29%2Fln%28278%29

OpenStudy (anonymous):

thanks guys i have one more question if you have the time

OpenStudy (anonymous):

that should be correct my guess is there was a problem putting it in the calculator

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

@satellite73 why do you times it by 4

OpenStudy (anonymous):

you need parentheses, but since you are on a computer, type it in to wolfram as i did

OpenStudy (anonymous):

oh damn!

OpenStudy (anonymous):

because i made a mistake it is .468 http://www.wolframalpha.com/input/?i=ln%2837800%29%2F%284ln%28278%29%29

OpenStudy (anonymous):

thats what i got .468 rounded off to .47

OpenStudy (anonymous):

yes you are right i am assuming that it was what not something like \[27\times 8^{4x}=37,800\] right?

OpenStudy (anonymous):

yes thanks